Please show work I don't understand!!!
11! / 4!(11-4)!
Answer is 330

Answer by Alan3354(69443)

Please show work I don't understand!!!
11! / 4!(11-4)!
-------------
n! means n factorial, = n*(n-1)*(n-2)*(n-3)...1
---
11! / 4!(11-4)! = 11!/(4!*7!)
= 11*10*9*8*7*6*5*4*3*2*1/(4*3*2*1*7*6*5*4*3*2*1)
-----
Obviously, all terms from 7 down cancel.
---
= 11*10*9*8/(4*3*2*1)
= 7920/24
= 330
==============================
If you have a calculator with a ! button, it will do these
If you don't have a calculator like that, get one.